Nadal Corporation manufactures custom molds for use in the extrusion industry. The company allocates manufacturing overhead based on machine hours. Selected data for costs incurred for Job 532 are as​ follows:

Direct materials used

​ $5,000

Direct labor cost

​ $36,800

Predetermined manufacturing overhead rate based on machine hours

​ $17

Direct labor hours worked

​3,100

Machine hours used

290

What is the manufacturing cost of Job​ 532?

A.

​$41,800

B.

​$94,500

C.

​$46,730

D.

​$4,930
Answer #1

Total manufacturing overhead=Predetermined overhead rate*Machine hours used

=(17*290)=$4930

Total manufacturing cost=Direct materials used+Direct labor cost+Total manufacturing overhead

=(5000+36800+4930)

=$46730.